OpenStudy (anonymous):

Solve 5x2 = –45x. Solve x2 = 14x – 49. Solve x2 + 18 = –9x. Solve 4x2 = 12x + 40.

OpenStudy (pfenn1):

Have you tried any of these? Is there one in particular that you are having trouble with?

OpenStudy (pfenn1):

The first equation, you can factor 5x and then rearrange and solve for x.,

OpenStudy (pfenn1):

The second equation can be rearranged into the form \[ax^2+bx+c=0\] Then the expression on the left can be factored into two expressions (x+A)(x+B) where (A + B) = -14 and AB = 49.

OpenStudy (pfenn1):

The third equation can be handled the same way but in this case (A + B) = 9 and AB = 18. This is the general idea.